Graham, a physics enthusiast, is asked by his instructor to visit a nearby carpentry. He observed there a hand–held powered reciprocating saw, say a sabre saw, which performs SHM in a vertical plane with an angular frequency of 350 rad s−1 . Assume that the tip of the toothed blade of the saw is at x = 0 with positive velocity when t = 0 and the amplitude of SHM is 2.40 cm. Write an equation, as a function of time, which shows the position of the tip of the toothed blade. Also calculate that how much will the blade take to travel from x = 0 to x = 1.20 cm? To 2.40 cm?
1
Expert's answer
2019-10-30T12:32:17-0400

The position of the tip of the toothed blade:


"x(t)=A\cos{\omega t}"

"x(t)=0.024\cos{350 t}"

Time from x = 0 to x = 2.40 cm:


"\frac{T}{4}=\frac{2\pi}{4\omega}=\frac{\pi}{2(350)}=0.00449\ s=4.49\ ms"

Time from x = 0 to x = 1.20 cm:


"0.012=0.024\cos{350 t}"

"t=0.00299\ s=2.99\ ms"